# ATLAS [Uncaptioned image]: Adaptive Transfer Scaling Laws for Multilingual Pretraining, Finetuning, and Decoding the Curse of Multilinguality

### v1

@@ 2 Experimental Setup
Datasets and Evaluation We use the MADLAD-400 dataset (Kudugunta et al., 2024), a popular CommonCrawl-based dataset with the most expansive coverage of languages, totaling over 400. The MADLAD-400 authors prioritized multilingual-specific curation for this pretraining corpus, including language detection, filtering, and preprocessing. For 50 languages, chosen to represent a range of language families, scripts, and degrees of resourcefulness, we partition a random test set, to evaluate vocabulary-insensitive loss (Tao et al., 2024) for fairer cross-lingual comparisons. We evaluate the fit of our fitted scaling laws using $R^{2}$ calculated on held-out test sets partitioned to assess specific dimensions of generalization: $R^{2}(N)$ for the largest model sizes, $R^{2}(D)$ for the largest token ranges, $R^{2}(C)$ for the largest compute runs, and $R^{2}(M)$ for unique training language mixtures not seen at training. We separate these dimensions in response to prior work that has demonstrated the importance of rigorous test-set splits in scaling laws research (Li et al., 2025a).
Model Training We pretrain models from 10M to 8B parameters, using hyperparameter choices similar to those trained in Kudugunta et al. (2024) (with updates as prescribed by recent work). In this work, we train both monolingual and multilingual models, varying the model size $N$, training tokens $D$, and multilingual data proportions. In particular, we focus on training monolingual models, bilingual models, and massively multilingual Unimax models (Chung et al., 2023). most experiments center around scale, token, and mixture variations on these languages: English, French, Chinese, Hindi, Swahili, Russian, Spanish, Portuguese, Japanese, though certain experiments evaluate across 50 languages. We use a 64k Sentence Piece Model vocabulary (Kudo, 2018). Across experiments we pretrain 280 monolingual models, 240 bilingual models, 120 multilingual mixtures, and we finetune 130 monolingual models from Unimax checkpoints, totaling over 750 independent training runs. Full experimental details, including the specific language choices, hyperparameters, and training mixtures are provided in full in Appendix˜B.
@@ 3 Adaptive Scaling Laws for Monolingual & Multilingual Settings
Challenges with existing scaling laws for multilingual modeling. In this section, we discuss our attempts to precisely model different monolingual and multilingual pretraining experiments. Scaling laws for English are typically based on Chinchilla (Hoffmann et al., 2022)—which we refer to as the Chinchilla Scaling Law (CSL). But beyond English, languages may have different scaling laws (Arnett & Bergen, 2025), and language resources are often limited, requiring a Data-Constrained Scaling Law (DCSL) (Muennighoff et al., 2024) to account for diminishing returns after multiple epochs. However, DCSL requires ample data both before and after one epoch to accommodate its two-stage fitting process. For high-resource languages (English, French, Chinese), it is costly to collect ample data beyond one epoch, and for low/mid-resource languages (even Hindi, Hebrew, and Swahili) it can be very difficult to collect sufficient observations below one epoch. [^1] In MADLAD-400 these languages have $0.3\%$ English tokens. At standard batch sizes, even frequently evaluating (every $1k$) isn’t enough to collect many (sometimes any) observations before 1 epoch is reached. As such, even for monolingual modeling, we require a scaling law that is data repetition-aware, and robust to different collection allowances for $(N,D)$.
@@
$\mathcal{D}_{\mathrm{eff}}=\underbrace{{\color[rgb]{0.1015625,0.38671875,0.6015625}\definecolor[named]{pgfstrokecolor}{rgb}{0.1015625,0.38671875,0.6015625}\mathcal{S}_{\lambda}\!\left(D_{t};U_{t}\right)}}_{\text{Monolingual}}\;+\;\underbrace{{\color[rgb]{0.1640625,0.5625,0.1640625}\definecolor[named]{pgfstrokecolor}{rgb}{0.1640625,0.5625,0.1640625}\sum_{i\in\mathcal{K}}\tau_{i}\,\mathcal{S}_{\lambda}\!\left(D_{i};U_{i}\right)}}_{\text{Transfer Languages}}\;+\;\underbrace{{\color[rgb]{0.7265625,0.1328125,0.140625}\definecolor[named]{pgfstrokecolor}{rgb}{0.7265625,0.1328125,0.140625}\tau_{\mathrm{other}}\,\mathcal{S}_{\lambda}\!\left(D_{\mathrm{other}};U_{\mathrm{other}}\right)}}_{\text{Other Languages}}$ (2)
@@ 5 The Curse of Multilinguality—Model Capacity Constraints
Figure 4: We empirically measure the relative degradation in loss (y-axis), as compared to a monolingual model of the same $(N,D)$, from adding pretraining languages (z-axis). Left: We fix $D=25B$ tokens, and vary the model size NN. Right: We fix $N=2B$ parameters, and vary the tokens $D$. The points are real empirical observations, averaged across languages. The mesh-grid is a surface estimate, using a cubic spline. Target language loss is most affected by the number of training languages, but this loss penalty declines for larger models with more capacity.
@@ 6 Pretrain or Finetune?
Figure 7: In terms of NN, we model the compute budget $C$ required for a model pretrained from scratch to outperform the model finetuned from the generic Unimax multilingual checkpoint. We estimate this relationship as $\text{log}(C)=10283128\,\times\,N^{1.65}$.
@@ 7 Acknowledgements
We thank Luke Zettlemoyer, Catherine Arnett and Stella Biderman for helpful discussions on the paper. We thank Biao Zhang and Xavier Garcia for the technical disccusions and feedback on early directions.
@@ Appendix A Extended Related Work
Scaling laws are used to study the behavior of deep learning systems on scaling training data and/or compute. Various researchers have observed scaling properties of generalization error with training data size and model capacity (Banko & Brill, 2001; Hestness et al., 2017; Amodei et al., 2016). Specifically, in the context of LLMs, Kaplan et al. (2020) explicitly propose a power law for the relationship between the loss $L$, number of parameters in the language model $N$, and the number of dataset tokens $D$. Later, Hoffmann et al. (2022) proposed a different formula:
@@
We build upon this form throughout the paper to fit our scaling laws. Other researchers, too, have built upon these scaling laws to study various aspects of scaling, such as neural network architectures (Clark et al., 2022; Tay et al., 2022; Frantar et al., 2023; Gu & Dao, 2023; Scao et al., 2022) or transfer learning (Henighan et al., 2020). Researchers have also investigated scaling properties in different domains of deep learning such as neural machine translation (Ghorbani et al., 2021; Gordon et al., 2021), vision language models like CLIP (Cherti et al., 2023; Henighan et al., 2020), vision (Alabdulmohsin et al., 2022; Zhai et al., 2022), reinforcement learning (Hilton et al., 2023; Jones, 2021; Gao et al., 2023), and recommendation systems (Ardalani et al., 2022).
@@ B.2 Experiment Design
Optimizer AdamW Commonly used in scaling law experiments (Loshchilov & Hutter, 2019; Hoffmann et al., 2022)

### v2

@@ 2 Experimental Setup
Datasets and Evaluation We use the MADLAD-400 dataset (Kudugunta et al., 2024), a popular CommonCrawl-based dataset with the most expansive coverage of languages, totaling over 400. The MADLAD-400 authors prioritized multilingual-specific curation for this pretraining corpus, including language detection, filtering, and preprocessing. For 50 languages, chosen to represent a range of language families, scripts, and degrees of resourcefulness, we partition a random test set, to evaluate vocabulary-insensitive loss (Tao et al., 2024) for fairer cross-lingual comparisons. Following (Muennighoff et al., 2024; He et al., 2024), we evaluate the fit of our parametric scaling laws using $R^{2}$ calculated on held-out test sets partitioned to assess specific dimensions of generalization: $R^{2}(N)$ for the largest model sizes, $R^{2}(D)$ for the largest token ranges, $R^{2}(C)$ for the largest compute runs, and $R^{2}(M)$ for unique training language mixtures not seen at training (more details in Subsection˜B.3). We separate these dimensions in response to prior work that has demonstrated the importance of rigorous test-set splits in scaling laws research (Li et al., 2025a).
Model Training We pretrain models from 10M to 8B parameters, using hyperparameter choices similar to those trained in Kudugunta et al. (2024) (with updates as prescribed by recent work). In this work, we train both monolingual and multilingual models, varying the model size $N$, training tokens $D$, multilingual data proportions, and the total training compute $C$. In particular, we focus on training monolingual models, bilingual models, and massively multilingual Unimax models (Chung et al., 2023). most experiments center around scale, token, and mixture variations on these languages: English, French, Chinese, Hindi, Swahili, Russian, Spanish, Portuguese, Japanese, though certain experiments evaluate across 50 languages. We use a 64k Sentence Piece Model vocabulary (Kudo, 2018). Across experiments we pretrain 280 monolingual models, 240 bilingual models, 120 multilingual mixtures, and we finetune 130 monolingual models from Unimax checkpoints, totaling over 750 independent training runs. Full experimental details, including the specific language choices, hyperparameters, and training mixtures are provided in full in Appendix˜B.
@@ 3 Adaptive Scaling Laws for Monolingual & Multilingual Settings
Challenges with existing scaling laws for multilingual modeling. In this section, we discuss our attempts to precisely model different monolingual and multilingual pretraining experiments. Scaling laws for English are typically based on Chinchilla (Hoffmann et al., 2022)—which we refer to as the Chinchilla Scaling Law (CSL). Hoffmann et al. (2022)’s two power-law terms (Equation˜1), allows us to separate how loss scales model size $N$ and data size $D$. $E$ is the irreducible loss (representing the entropy of natural text), ($A$, $B$) are learned coefficients and and ($\alpha$, $\beta$) are scaling exponents for for model size and data size respectively. Beyond English, languages may have different scaling laws (Arnett and Bergen, 2025), and language resources are often limited, requiring a Data-Constrained Scaling Law (DCSL) (Muennighoff et al., 2024) to account for diminishing returns after multiple epochs. However, DCSL requires ample data both before and after one epoch to accommodate its two-stage fitting process. For high-resource languages (English, French, Chinese), it is costly to collect ample data beyond one epoch, and for low/mid-resource languages (even Hindi, Hebrew, and Swahili) it can be very difficult to collect sufficient observations below one epoch. [^1] In MADLAD-400 these languages have $0.3\%$ English tokens. At standard batch sizes, even frequently evaluating (every $1k$) isn’t enough to collect many (sometimes any) observations before 1 epoch is reached. As such, even for monolingual modeling, we require a scaling law that is data repetition-aware, and robust to different collection allowances for $(N,D)$.
@@
$\mathcal{D}_{\mathrm{eff}}=\underbrace{{\color[rgb]{0.1015625,0.38671875,0.6015625}\definecolor[named]{pgfstrokecolor}{rgb}{0.1015625,0.38671875,0.6015625}\mathcal{S}_{\lambda_{t}}\!\left(D_{t};U_{t}\right)}}_{\text{Monolingual}}\;+\;\underbrace{{\color[rgb]{0.1640625,0.5625,0.1640625}\definecolor[named]{pgfstrokecolor}{rgb}{0.1640625,0.5625,0.1640625}\sum_{i\in\mathcal{K}}\tau_{i}\,\mathcal{S}_{\lambda_{i}}\!\left(D_{i};U_{i}\right)}}_{\text{Transfer Languages}}\;+\;\underbrace{{\color[rgb]{0.7265625,0.1328125,0.140625}\definecolor[named]{pgfstrokecolor}{rgb}{0.7265625,0.1328125,0.140625}\tau_{\mathrm{other}}\,\mathcal{S}_{\lambda_{\mathrm{other}}}\!\left(D_{\mathrm{other}};U_{\mathrm{other}}\right)}}_{\text{Other Languages}}$ (2)
@@ 5 The Curse of Multilinguality—Model Capacity Constraints
Figure 4: We empirically measure the relative degradation in loss (y-axis), as compared to a monolingual model of the same $(N,D)$, from adding pretraining languages (z-axis). Left: We fix $D=25B$ tokens, and vary the model size $N$. Right: We fix $N=2B$ parameters, and vary the tokens $D$. The points are real empirical observations, averaged across languages. The mesh-grid is a surface estimate, using a cubic spline. Target language loss is most affected by the number of training languages, but this loss penalty declines for larger models with more capacity.
@@ 6 Pretrain or Finetune?
Figure 7: In terms of NN, we model the compute budget $C$ required for a model pretrained from scratch to outperform the model finetuned from the generic Unimax multilingual checkpoint. We estimate this relationship as $C=10283128\,\times\,N^{1.65}$.
@@
7 Conclusion
In this work, we answer several scaling questions on training multilingual models, validating our findings on scales up to 8B model parameters. First, we introduce a novel scaling law formulation called Adaptive Transfer Scaling Law (ATLAS) that significantly improves scaling law fit by more than 0.3 $R^{2}$ for multilingual mixtures. To aid our analysis, we also derive transfer scores for 38 langauages. We find that English is often a strong positive transfer language, alongside languages with shared script and language families, though we note that this transfer is often asymmetric. However, there is still a compute-efficiency tax when we train multilingual models. We quantify this "curse of multilinguality" and provide a mathematical trade-off between model capacity and language coverage: we find that the performance penalty from adding languages is mitigated more effectively by scaling model size ($N$) rather than training tokens ($D$). Finally, we quantitatively answer a critical practical question for practitioners: determining the compute crossover point between pretraining and finetuning. We find that depending on the language, for a budget under 144-283B tokens, finetuning a general-purpose multilingual checkpoint is more efficient. Collectively, these contributions provide valuable guidelines for practitioners determining data mixtures, language coverage and computational requirements for creating multilingual foundation models.
@@ 8 Acknowledgements
We thank Luke Zettlemoyer, Catherine Arnett and Stella Biderman for helpful discussions on the paper. We thank Biao Zhang and Xavier Garcia for the technical discussions and feedback on early directions.
@@ Appendix A Extended Related Work
Scaling laws are used to study the behavior of deep learning systems on scaling training data and/or compute. Various researchers have observed scaling properties of generalization error with training data size and model capacity (Banko and Brill, 2001; Hestness et al., 2017; Amodei et al., 2016). Specifically, in the context of LLMs, Kaplan et al. (2020) explicitly propose a power law for the relationship between the loss $L$, number of parameters in the language model $N$, and the number of dataset tokens $D$. Later, Hoffmann et al. (2022) proposed a different formula:
@@
We build upon this form throughout the paper to fit our scaling laws. Other researchers, too, have built upon these scaling laws to study various aspects of scaling, such as neural network architectures (Clark et al., 2022; Tay et al., 2022; Frantar et al., 2023; Gu and Dao, 2023; Scao et al., 2022) or transfer learning (Henighan et al., 2020). Researchers have also investigated scaling properties in different domains of deep learning such as neural machine translation (Ghorbani et al., 2021; Gordon et al., 2021), vision language models like CLIP (Cherti et al., 2023; Henighan et al., 2020), vision (Alabdulmohsin et al., 2022; Zhai et al., 2022), reinforcement learning (Hilton et al., 2023; Jones, 2021; Gao et al., 2023), and recommendation systems (Ardalani et al., 2022).
@@ B.2 Experiment Design
Optimizer AdamW Commonly used in scaling law experiments (Loshchilov and Hutter, 2019; Hoffmann et al., 2022)